# Characterizing Translation-Invariant Bell Inequalities using Tropical Algebra and Graph Polytopes

### v1

@@ Acknowledgements.
We thank Albert Aloy and Zizhu Wang for sharing their insights into the Bell polytope, and Jinfu Chen for feedback on an early version of the manuscript. T.S. would like to thank Stijn Cambie for pointing out relevant references in the graph theory literature. P.E. and J.T. acknowledge the support received by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. P.E. additionally acknowledges the support received through the NWO-Quantum Technology programme (Grant No. NGF.1623.23.006). J.T. and E.V. acknowledge the support received from the European Union’s Horizon Europe research and innovation programme through the ERC StG FINE-TEA-SQUAD (Grant No. 101040729). T.S. is supported by Research foundation – Flanders (FWO) - Grant Number 1219723N. This publication is part of the ’Quantum Inspire - the Dutch Quantum Computer in the Cloud’ project (with project number [NWA.1292.19.194]) of the NWA research program ’Research on Routes by Consortia (ORC)’, which is funded by the Netherlands Organization for Scientific Research (NWO). The views and opinions expressed here are solely those of the authors and do not necessarily reflect those of the funding institutions. Neither of the funding institutions can be held responsible for them.
@@ B.2.1 Construction of ${\operatorname{err}_{\Gamma}(c,r)}$
We can find a linear function $\xi$ such that the unique minimum of $\xi$ on $P_{N}(\Gamma)$ is achieved at $p$, and the unique minimum of $\xi$ on $N\cdot p_{*}(\Gamma)$ is achieved at $w(c)$ for some $c$ in $C_{F}$. Indeed: start with any $\xi$ that has its unique minimum on $P_{N}(\Gamma)$ at $p$ (this exists because $p$ is a vertex), perturb the coefficients such that all numbers $\xi(w(c))$ are distinct, then add to it a linear function that is 00 on $F$ and takes large positive values on the $w(c)$ with $c\notin C_{F}$.
@@ B.2.2 Proof of finiteness and inefficient computation
Now we need to take the set of all $c$-irreducible lattice points in $C_{c,r}$. Clearly whenever all $\lambda_{c^{\prime}}$ are integers, Equation 89 is a lattice point. With some more work, one sees that actually every lattice point in $C_{c,r}$ can be written as Equation 89 with all coefficients integers 666the idea is to add a large multiple of $W(C_{0})$ so all entries become positive, and then use decomposition of a cycle into simple cycles. The $c$-irreducibility only depends on which of the $\lambda_{c^{\prime}}$ (with $c^{\prime}\neq c$) are zero and which are not.
@@ B.2.3 Making the computation more efficient
$\bullet$ a simple cycle $c_{0}\in\mathcal{C}_{\Gamma}$
$\bullet$ an interger $r\in\{0,\dots,\ell_{c_{0}}-1\}$
@@
[…]
$\lambda_{c_{0}}\leftarrow-\lfloor\frac{1}{\ell_{c_{0}}}\sum_{c\in S}\lambda_{c}\ell_{c}\rfloor$LCommentLattice point in the cone $C_{c_{0},r}$
[…]
$\epsilon$_in_$\mathcal{E}_{c_{0},r}^{\prime}$leftarrow False
[…]
@@ B.3 The case of $K_{4}$: results and example
cycle length ##\## cycles $r=0$ $r=1$ $r=2$ $r=3$
@@
cycle length ##\## cycles r=0 r=1 r=2 r=3
@@ Appendix C Symmetry classes for TI-Bell inequalities
In general, we consider that two inequalities are equivalent if they are equal under the permutation of parties, input and outcome labels. For one party, there are $12m-$ 1 generators of permutations. The first $m-1$ generators are the input labels permutations $\alpha_{x}\mapsto\alpha_{\sigma(x)}$, where $\sigma$ is a generator of the symmetric group $\text{Sym}(m)$. The $m$ remaining generators are given by the outcome labels permutations for input $x^{\prime}$: $\alpha_{x}\mapsto-\alpha_{x}$ if $x=x^{\prime}$. By composing these $12m-$ 1 generators, one gets all possible input and outcome label permutations for a single party. Since we are working with TI inequalities, the same input and outcome permutations must be applied to all the parties. Finally, in TI inequalities, there is also one possible party permutation. It corresponds to the flipped permutation: $ABCD\rightarrow DCBA$. Under this permutation, the neighbors of a given party remain unchanged. This holds for any interaction range $R$. We have thus in total $2m$ generators for TI-$R$ inequalities.
@@
For example in the case $R=m=2$, there are four permutation generators: parties are flipped $\alpha_{x,y}^{i,j}\rightarrow\alpha_{y,x}^{i,j}$, inputs are flipped, outcomes are flipped when input is 00 and outcomes are flipped when input is $1$:
@@
where all the input indices are taken modulo $m$. In the $(N,2)$ scenario with $R=2$ and $N$ divisible enough, the projected polytope has $32372$ facets and there are $2102$ different symmetric classes  [13].
@@ Appendix D Renormalization computation
Since $\mathcal{P}_{F^{\prime}}$ is defined by homogeneous linear inequalities it is a so-called polyhedral cone. A version of the Minkowski-Weyl theorem 777to be precise, we need the additional requirement that the cone is pointed, i.e. that it doesn’t contain a linear space. This follows from the injectivity assumption. states that such a cone is the conical hull of finitely many vectors $\bm{a}_{1},\dots,\bm{a}_{c}$:

### v2

@@ Acknowledgements.
We thank Albert Aloy and Zizhu Wang for sharing their insights into the Bell polytope, and Jinfu Chen for feedback on an early version of the manuscript. T.S. would like to thank Stijn Cambie for pointing out relevant references in the graph theory literature. F.M. would like to thank Marianne Johnson and Henry Wynn for helpful discussions on tropical matrix algebra. F.M. was partially supported by the Research Foundation – Flanders (FWO), Grant Numbers G0F5921N (Odysseus, Type I) and G023721N, as well as the grant iBOF/23/064 from KU Leuven. These grants also supported the research visits of M.H. from Leiden to Leuven, which contributed to this work. P.E. and J.T. acknowledge the support received by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. P.E. additionally acknowledges the support received through the NWO-Quantum Technology programme (Grant No. NGF.1623.23.006). J.T. and E.V. acknowledge the support received from the European Union’s Horizon Europe research and innovation programme through the ERC StG FINE-TEA-SQUAD (Grant No. 101040729). T.S. is supported by Research foundation – Flanders (FWO) - Grant Number 1219723N. This publication is part of the ’Quantum Inspire - the Dutch Quantum Computer in the Cloud’ project (with project number [NWA.1292.19.194]) of the NWA research program ’Research on Routes by Consortia (ORC)’, which is funded by the Netherlands Organization for Scientific Research (NWO). The views and opinions expressed here are solely those of the authors and do not necessarily reflect those of the funding institutions. Neither of the funding institutions can be held responsible for them.
Authors contribution
Conceptualization: T.S., F.M., J.T. Methodology: M.H., E.V., T.S., P.E., F.M., J.T. Software: M.H., E.V., T.S. Validation: T.S., P.E., J.T. Formal analysis: M.H., E.V., T.S., F.M. Data curation: M.H., E.V., P.E. Writing - Original Draft: M.H., E.V., T.S., F.M. Writing - Review & Editing: P.E., J.T. Visualization: M.H., E.V., P.E. Supervision: T.S., P.E., J.T. Project administration: J.T. Funding acquisition: T.S., P.E., F.M., J.T.
@@ B.2.1 Construction of $\operatorname{err}_{\Gamma}(c,r)$}
We can find a linear function $\xi$ such that the unique minimum of $\xi$ on $P_{N}(\Gamma)$ is achieved at $p$, and the unique minimum of $\xi$ on $N\cdot p_{*}(\Gamma)$ is achieved at $w(c)$ for some $c$ in $C_{F}$. Indeed: start with any $\xi$ that has its unique minimum on $P_{N}(\Gamma)$ at $p$ (this exists because $p$ is a vertex), perturb the coefficients such that all numbers $\xi(w(c))$ are distinct, then add to it a linear function that is 0 on $F$ and takes large positive values on the $w(c)$ with $c\notin C_{F}$.
@@ B.2.2 Proof of finiteness and inefficient computation
Now we need to take the set of all $c$-irreducible lattice points in $C_{c,r}$. Clearly whenever all $\lambda_{c^{\prime}}$ are integers, Equation 89 is a lattice point. With some more work, one sees that actually every lattice point in $C_{c,r}$ can be written as Equation 89 with all coefficients integers (the idea is to add a large multiple of $W(C_{0})$ so all entries become positive, and then use decomposition of a cycle into simple cycles). The $c$-irreducibility only depends on which of the $\lambda_{c^{\prime}}$ (with $c^{\prime}\neq c$) are zero and which are not.
@@ B.2.3 Making the computation more efficient
$\bullet$ a finite oriented graph $\Gamma$ with self-loops, but no parallel edges $\bullet$ a simple cycle $c_{0}\in\mathcal{C}_{\Gamma}$ $\bullet$ an interger $r\in\{0,\dots,\ell_{c_{0}}-1\}$
$\triangleright$  Construct the cone $C_{c_{0},r}$ with vertex $r\cdot w_{c_{0}}$ and rays $\{(w_{c}-w_{c_{0}})\mid c\in\mathcal{C}_{\Gamma}\}$   $\triangleleft$
$C_{c_{0},r}\leftarrow\textsc{cone}(r\cdot w_{c_{0}},\{w_{c}-w_{c_{0}}\}_{c\in\mathcal{C}_{\Gamma}})$
$\triangleright$  Compute the triangulation of the cone   $\triangleleft$
$\Delta\leftarrow\textsc{triangulate}(C_{c_{0},r})$
$\triangleright$  Discard simplices according to claims 2,3 and 4   $\triangleleft$
$\Delta\leftarrow\textsc{apply\_claims}(c_{0},\Delta)$
$\triangleright$  For all the remaining simplices, find the points that are in $\mathcal{E}^{\prime}_{c,r}$. Start with empty set   $\triangleleft$
$\mathcal{E}^{\prime}_{c_{0},r}\leftarrow\emptyset$
for $S\in\Delta$ do
   $\triangleright$  According to the remark 2, $\lambda_{c}=1,\dots,\operatorname{lcm}(\ell_{c_{0}},\ell_{c})/\ell_{c}$   $\triangleleft$
   for $(\lambda_{c})_{c\in S}$ with $\lambda_{c}=1,\dots,\frac{\operatorname{lcm}(\ell_{c_{0}},\ell_{c})}{\ell_{c}}+1$ do
    $\triangleright$  Since we want points in the cone, we need $\lambda_{c_{0}}\ell_{c_{0}}+\sum_{c\in S}\lambda_{c}\ell_{c}=r$   $\triangleleft$
    if $\sum_{c\in S}\lambda_{c}\ell_{c}\mod\ell_{c_{0}}=r$ then
      $\triangleright$  Since we want lattice points, $\lambda_{c_{0}}$ is the integer satisfying the cone constraint   $\triangleleft$
      $\lambda_{c_{0}}\leftarrow-\lfloor\frac{1}{\ell_{c_{0}}}\sum_{c\in S}\lambda_{c}\ell_{c}\rfloor$
      $\triangleright$  Lattice point in the cone $C_{c_{0},r}$   $\triangleleft$
      $\epsilon\leftarrow\lambda_{c_{0}}W_{c_{0}}+\sum_{c\in S}\lambda_{c}W_{c}$
      if $\epsilon$ is $c_{0}$-irreducible then
       $\triangleright$  $\epsilon\in\mathcal{E}_{c_{0},r}$ because it is a $c_{0}$-irreducible lattice point in the cone. Now we check if $\epsilon$ is also in $\mathcal{E}_{c_{0},r}^{\prime}$   $\triangleleft$
@@
       for $c\in S$ do
         $\epsilon^{\prime}\leftarrow\epsilon+\operatorname{lcm}(\ell_{c_{0}},\ell_{c})(w_{c_{0}}-w_{c})$
         if $\epsilon^{\prime}$ is $c_{0}$-irreducible and has no negative entries outside $c_{0}$ then
          $\triangleright$  $\epsilon^{\prime}\in\mathcal{E}_{c_{0},r}$ because it is a $c_{0}$-irreducible lattice point in the cone. Thus $\epsilon\notin\mathcal{E}_{c_{0},r}^{\prime}$.   $\triangleleft$
@@
         $\mathcal{E}^{\prime}_{c_{0},r}\leftarrow\mathcal{E}^{\prime}_{c_{0},r}\cup\{\epsilon\}$
$\triangleright$  Construct the polyhedron $\tilde{C}_{c_{0},r}$ with the points $\mathcal{E}^{\prime}_{c_{0},r}$ and rays $\{(w_{c}-w_{c_{0}})\mid c\in\mathcal{C}_{\Gamma}\}$   $\triangleleft$
$\tilde{C}_{c_{0},r}\leftarrow\textsc{cone}(\mathcal{E}^{\prime}_{c_{0},r}\;,\{w_{c}-w_{c_{0}}\}_{c\in\mathcal{C}_{\Gamma}})$
$\triangleright$  The error list correspond to the vertices of $\tilde{C}_{c_{0},r}$   $\triangleleft$
$\widetilde{\operatorname{err}}_{\Gamma}(c_{0},r)\leftarrow\textsc{vertices}(\tilde{C}_{c_{0},r})$
@@ B.3 The case of $K_{4}$: results and example
cycle length #\# cycles $r=0$ $r=1$ $r=2$ $r=3$
@@
cycle length #\# cycles r=0 r=1 r=2 r=3
@@ Appendix C Symmetry classes for TI-Bell inequalities
In general, we consider that two inequalities are equivalent if they are equal under the permutation of parties, input and outcome labels. For one party, there are 2m12m-1 generators of permutations. The first m1m-1 generators are the input labels permutations $\alpha_{x}\mapsto\alpha_{\sigma(x)}$, where $\sigma$ is a generator of the symmetric group $\text{Sym}(m)$. The $m$ remaining generators are given by the outcome labels permutations for input $x^{\prime}$: $\alpha_{x}\mapsto-\alpha_{x}$ if $x=x^{\prime}$. By composing these 2m12m-1 generators, one gets all possible input and outcome label permutations for a single party. Since we are working with TI inequalities, the same input and outcome permutations must be applied to all the parties. Finally, in TI inequalities, there is also one possible party permutation. It corresponds to the flipped permutation: $ABCD\rightarrow DCBA$. Under this permutation, the neighbors of a given party remain unchanged. This holds for any interaction range $R$. We have thus in total $2m$ generators for TI-$R$ inequalities.
@@
For example in the case $R=m=2$, there are four permutation generators: parties are flipped $\alpha_{x,y}^{i,j}\rightarrow\alpha_{y,x}^{i,j}$, inputs are flipped, outcomes are flipped when input is 0 and outcomes are flipped when input is $1$:
@@
where all the input indices are taken modulo $m$. In the $(N,2)$ scenario with $R=2$ and NN divisible enough, the projected polytope has $32372$ facets and there are $2102$ different symmetric classes  [13].
@@ Appendix D Renormalization computation
Since $\mathcal{P}_{F^{\prime}}$ is defined by homogeneous linear inequalities it is a so-called polyhedral cone. A version of the Minkowski-Weyl theorem (to be precise, we need the additional requirement that the cone is pointed, i.e. that it doesn’t contain a linear space, this follows from the injectivity assumption) states that such a cone is the conical hull of finitely many vectors $\bm{a}_{1},\dots,\bm{a}_{c}$: